From the function, f(x) = (3x ^ 2 - 8x - 3) / (2x ^ 2 + 7x - 4) ,




a. Construct table of values




b. Find the following:




vertical and horizontal asymptotes




x - ir intercepts or zeroes upstingit)




y- intercept,

1
Expert's answer
2021-11-23T15:12:23-0500

X=0 f(x)=3/4

X=1 f(x)=-8/5

X=-1 f(x)=-8/9

X=2 f(x)=-7/18


b. vertical asymptotes:

"2x^2+7x-4=0"

"D=49+32=81"

"x_1=(-7+9)\/4=1\/2"

"x_2=(-7-9)\/4=-4"

Vertical asymptotes are x=1/2, x=-4


Horizontal asymptotes"f(x)=3x^2\/2x^2=3\/2"


Y-interceptiins: X=0, f(x)=3/4

X-interceptions: f(x)=0, "3x^2-8x-3=0"

"D=64+36=100"

"x_1=(8+10)\/6=3"

"x_2=(8-10)\/6=-1\/3"

X-interceptions: x=3, f(x)=0

x=-1/3, f(x)=0